# A Tale of Two Imperatives: Privacy and Explainability

### v3

@@ 4 Localization Assumption & Privacy Invariance Score
Right-to-Privacy (RTP) Thomson (1975) and Right-to-Explanation (RTE) Vredenburgh (2022) are two inalienable aspects of modern-day ML software. We have already mentioned that previous works have shown potential security breaches leveraging the explanation in section 3.1. So, RTE can hamper RTP but is the converse true? If yes, How can we quantify the same? We first propose the desideratum for explanations in this setting and then run extensive experiments to quantitatively judge the explainers.
@@ 4.1 Description of the setting
It is shown that adversaries can leverage sensitive information from explanations but due to post-processing of DP, any DP model will always output private explanation Dwork et al. (2014). We want to check the extent to which the explanations from a DP model can be used as a proxy for that of the non-private model here. Formally, for a non-private model $\mathcal{M}$, its private counterpart $\mathcal{M^{\prime}}$; consider a set of post-hoc explainers $\mathcal{E}$ used for auditing $\mathcal{M}$ and $\mathcal{M^{\prime}}$. For an input $x\in X$ (where $X$ is the valid input space), each $\mathcal{I}\in\mathcal{E}$ produces explanations $s$ and $s^{\prime}$ of $x$ for $\mathcal{M}(x)$ and $\mathcal{M^{\prime}}(x)$ respectively, Privacy Invariance Score (PIS) over a given tuple $(\mathcal{M},\mathcal{M^{\prime}},x,\mathcal{I})$ is defined as follows.
Definition 1. Privacy Invariance Score (PIS): Given a tuple $(\mathcal{M},\mathcal{M^{\prime}},x,\mathcal{I})$ given $\mathcal{M}(x)$ = $\mathcal{M^{\prime}}(x)$, PIS is defined as $sim(s,s^{\prime})$. Where $sim(\cdot,\cdot)$ is a similarity metric, $s=\mathcal{I}(\mathcal{M},x)$, and $s^{\prime}=\mathcal{I}(\mathcal{M^{\prime}},x)$.
@@ 4.2 Localization Assumption
As discussed earlier, different works evaluated explainers in various ways. These ad-hoc norms are often unique to each paper and inconsistent Jacovi and Goldberg (2020). However, practitioners have proposed some necessary axiomatic desiderata Lyu et al. (2024) such as the well-established Implementation Invariance (II) criterion proposed by Sundararajan et al. Sundararajan et al. (2017).
According to II if $\mathcal{I}$ is faithful and $\mathcal{M}(x)=\mathcal{M^{\prime}}(x)$ $\forall x\in X$, i.e. $\mathcal{M,M^{\prime}}$ are functionally equivalent (FE) then $s=s^{\prime}$ $\forall x$. Commonly used explainers like Integrated Gradient Sundararajan et al. (2017), SmoothGrad Smilkov et al. (2017), DeepLift Li et al. (2021), layerwise relevance propagation (LRP) Montavon et al. (2019) etc had been extensively accessed based on II Sundararajan et al. (2017).
Following the line of prior explainability research, Jacovi et al. Jacovi and Goldberg (2020) formally proposed The Model Assumption (MA):
@@
Nevertheless, in our setting, as the accuracy of $\mathcal{M^{\prime}}$ is generally less Ji et al. (2014), and we cannot say $\mathcal{M}$ and $\mathcal{M^{\prime}}$ are FE Sundararajan et al. (2017). Furthermore, as $X$ in practice can be arbitrarily large (theoretically could be countably infinite), quantifying whether an explainer ($\mathcal{I}$) is sufficiently trustworthy or not is often impractical. Also, finding a counterexample that violates the condition implying $\mathcal{I}$ is not faithful is computationally expensive. Hence, we modify MA and propose the localization assumption for evaluating the explainers’ quality in our setting.
@@
    Alignment with the original model (Alignment): The proxy model must closely align with the original model in its predictions. In our study, we mesaure the same with the agreement on the “hard predictions” obtained from $\mathcal{M}$ and $\mathcal{M^{\prime}}$ over the test set. (Dis)agreement has been extensively used in previous studies due to its its simplicity and interpretability Shamir and Coviello (2020); Skalak et al. (1996); Liu et al. (2022); Milani Fard et al. (2016).
@@ 4.3 The Notion of Similarity
As previously discussed in PIS, we aim to measure the similarity between pair of explanations ; it is essential to account for two key factors here: the context of comparison and human understanding of that comparison. In this framework, given $(\mathcal{M},\mathcal{M^{\prime}},x,\mathcal{I})$ and $\mathcal{M}(x)=\mathcal{M^{\prime}}(x)$, we seek to investigate two primary aspects:
@@
Firstly, explanations generally contain both positive and negative attributions [^4] But a few explainers such as Saliency only outputs positive attributions; Saliency in particular, outputs the absolute value of the attributions.. For a classification, the positive attributions contribute in the favor of the classification while the negative ones don’t. Keeping this in mind, in order to address the first question, we start by calculating the number of places in both explanations where they yield different types of attribution scores; we denote this as the disagreement score (DS, measured in %). As both s and $s^{\prime}$ comprise attribution scores (per pixel or per element) that either positively contribute to the classification or do not, DS serves as a (crude) measure for assessing how much the reasoning of the models, as indicated by $\mathcal{I}$, diverges at the pixel or element level. Consequently, in a complementary manner, we focus on identifying the set of pixels selected by both models for classification. For instance, if the disagreement score is 00, it indicates that both models select the exact same set of positive and negative pixels (or elements) for classification. Conversely, a certain percentage of disagreement implies that $\mathcal{I}$ reveals differences of reasoning in how $\mathcal{M}$ and $\mathcal{M^{\prime}}$ interpret pixel(s) for the same classification task.
From a human evaluation perspective though, analysis on DS is context-sensitive. While evaluating DS may not be crucial in all applications, it becomes critically important in high-stakes domains, such as biomedical software, where discrepancies could lead to significant confusion. In these contexts, both position and attribution of information are vital for accurate treatment decisions. To address this, our system evaluates disagreement at an atomic level, ensuring precision and thoroughly scrutinizing the explainer when a specific threshold of disagreement is identified for an image. In our study, we set the DS threshold 15%. We employ DS primarily as a sanity check in our experiment for the comparability of model explanations.
To tackle the second question, after measuring DS and eliminating the ones not adhering the threshold, we assess how the attribution assignments correlate with each other. However, the correlation at pixel-level attributions is inherently a crude estimate. Given that each input image includes over 50,000 pixels, computing correlation directly is computationally demanding. Thus, we employ a two-step approach here.
Firstly, we segregate the positive attributions from the negative ones. This distinction is crucial because positive pixels contribute positively to the classification, while negative attribution scores may indicate a range of scenarios: a pixel could primarily (or completely) contribute to other classes, contribute negatively to the output class, or any combination of these two Ancona et al. (2017); Samek et al. (2017). This is why we solely focus on the pairwise positive pixels from both explanations and compute their rank correlation with the nonparametric Kendall Tau [^5] A natural question might arise: What if all the pixels have negative attributions? As we are taking the explanation against the “hard prediction”, if an explanation is not having any pixel with positive attributions yet it predicted the selected class; we’d argue its faithfulness. As a negative attribution across all pixels would suggest that the model made its prediction without relying on any relevant pixel/element from the input image. Moreover, throughout our experiments we never encountered such instance and positive attributed pixels were never $<60\%$. We empirically set a (crude) baseline (in a context aware fashion) of $0.5$ ($50\%$ positively correlated) for the comparison. Our approach is easily implementable with commonly used libraries such as tf-explain [^6] https://tf-explain.readthedocs.io/en/latest/ and Captum [^7] https://captum.ai/.
4.4 Other notions for Similarity
Can cosine similarity be an alternative? : Mostly No, as in such high dimensions the large random vectors are almost always almost orthogonal if they are not (condensely) clustered Mohammadi and Petridis (2022); Wyner (1967). Moreover, unlike the correlation and disagreement cosine is not that much nuanced as this doesn’t compare the Condorcet pairs but imposes a vague sense of how much the vectors are ‘aligned’ defined over their inner product space in such high dimension which doesn’t go with our objective.
Can $L_{P}$ norms be a better metric? No, we need to normalize before comparing the vectors, which kills one degree of freedom, which is not the case with correlation.
@@ 5.1 Description of the Dataset
Our dataset comprises 2,000 Pneumonia cases sourced from the Chest X-ray dataset by Sharma et al. Sharma (2020), and 2,000 TB cases randomly sampled from the NIAID TB Portal Program dataset National Institute of Allergy and Infectious Diseases (n.d.). To create the ‘Normal’ subset, we include an equal split of 1,000 unaffected Pneumonia samples from the uneffected class in the aforsaid Chest X-ray dataset Sharma (2020) and 1,000 unaffected TB samples from Rahman et al. Rahman et al. (2020), totaling 2,000 normal cases. For evaluation, the test set contains 200 images for each class.
@@
This paper, unlike prior studies, is not focused on the empirical studies of private models in healthcare Zerka et al. (2020); Naresh et al. (2023); Khalid et al. (2023). Instead, we concentrate on the applicability of commonly used explainers in a privacy preserving environment. Our aim is to investigate whether RTE and RTP can be achieved simultaniously. As a result, rather than a large dataset where extensive experimentation with DP under a closely controlled environment is computationally expensive and cumbersome Subramani et al. (2021), we focus on first making a dataset of an appropriate size where we have considered the commonly available and utilized disease class from the previous studies: Pneumonia, Tuberculosis (TB). However, after our primary experiment, we devise the same setup for another benchmark dataset: CIFAR-10 but we arrived at similar conclusion. Details can be found in Appendix B.
[…]
The determination of privacy requirements and expectations is a critical aspect of our experimental setup. However, selecting $\epsilon$ comes with a notable dichotomy:
The theoretical school advocates for stringent privacy guarantees, proposing $\epsilon\leq 1$. This conservative approach provides robust worst-case guarantees against inference attacks. For instance, in the case of, $\epsilon\leq 0.1$ ensures that no membership inference attack (MIA) can outperform random guessing by more than a marginal 2.5% (i.e., 52.5% success rate).
Conversely, industrial implementations often employ substantially larger values, typically $\epsilon>7$. According to the theoretical school such choices, while pragmatic, significantly increase vulnerability. With such choices of $\epsilon$, system becomes severely vulnerable with the worst case MIA attack succedding rate exceeding $99\%$.
This stark contrast stems from divergent underlying assumptions Lowy et al. (2024):
  • The theoretical model assumes an adversary with near-complete knowledge of the training set, lacking information on only a single data point.
  • It also presupposes uniform privacy guarantees across all data points.
Very recently, Lowy et al. (2024) has shown why often these assumptions are not relevant in practical and industrial settings and DP with large $\epsilon$ (even $\epsilon\geq 7$) can sufficiently defend against practical and implementable MIAs. Furthermore, in reality, training data is typically either fully public (subject to regulatory compliance) or entirely private. Particularly in several high-stakes domain, data breaches and unauthorized access to sensitive information (e.g., healthcare records) often result in comprehensive system restrictions, as exemplified by Italy’s 2023 temporary ban on ChatGPT due to privacy concerns McCallum (2023).
Nevertheless, this study is not about how to choose ‘$\epsilon$’ for the biomedical software(s). Moreover, such questions are open-ended and as far as healthcare is concerned, it is more on the side of the authoritative bodies on how they impose the guidelines rather than individual stakeholders. In this study, we follow the guidelines proposed by Google where they introduced a three-tier system Ponomareva et al. (2023).
  • Tier 1: $\epsilon\leq 1$
  • Tier 2: $\epsilon\leq 10$
  • Tier 3: $\epsilon>10$
Tier 1 matches the theoretical view but loses too much utility in real-world settings. Tier 2 is advocated by Google as it adheres to practical constraints and yields reasonable privacy-utility trade-off. Lastly, tier 3 is avoided in this study as it makes the system excessively vulnerable.
@@
Our primary investigation examines six discrete $\epsilon$ values: $0.4,0.7,1,4,7,$ and $10$. This range enables a comprehensive analysis of the privacy-utility trade-off spanning across tier 1 and tier 2.
@@ 5.3 Description of the Explainers
We primarily selected gradient-based explainers which are predominantly used in computer vision tasks Buhrmester et al. (2021) — specifically, Saliency, SmoothGrad, Integrated Gradients, Grad-Shap and, Grad-CAM. These gradient-based explainers rely on the sensitivity the model shows to the features subject to the output it obtains. In other words, these explainers leverage the gradient of the output (or any selected class of interest) w.r.t. the input features888it can either directly be the element(s) of the given input or the intermediate ones coming from a layer inside the model.. For example, Saliency computes the partial derivatives of the output with respect to the input given, Integrated Gradients computes the average gradient while the input varies along a linear path from a baseline. SmoothGrad involves adding noise to the input image and generating multiple saliency maps, each corresponding to a slightly different noisy version of the input image; the saliency maps are then averaged to produce a ‘smoothed’ saliency map. Grad-CAM also involves taking gradients of the target output with respect to the given layer. From perturbation-based explainers, we included only SHAP Lundberg and Lee (2017), specifically Grad-Shap, as core perturbation methods are less widely adopted for computer vision. Grad-Shap can be viewed as an approximation of Integrated Gradients by computing the expectations of gradients for different baselines [^9] Captum . Other perturbation-based methods like LIME Ribeiro et al. (2016) assign weights to super-pixels rather than individual pixels as a result, we cannot compare the same with other methods we chose which output per-pixel or per-element attribution values, Occlusion can lead to OOD artifacts and is not widely used Chang et al. (2018). Additionally, some other explainers such as DeepLIFT, LRP disregard even basic faithfulness tests Sundararajan et al. (2017).
@@ 5.4 Description of the Networks
Disease detection using chest X-ray is a well-regarded case study. In previous works Al-qaness et al. (2024), the choice of CNNs is significantly prominent and in this work also we have chosen the famous CNNs that had been used in previous research. Particularly, We selected ResNet-34 He et al. (2016), EfficientNet-v2101010For this study, we used the ‘small’ version from the EfficientNet-v2 family, referred to throughout the paper as EfficientNet-v2. Tan and Le (2021), and DenseNet-121 Huang et al. (2017) for our analysis due to their competitive performance. Importantly, we ensured that the DP counterparts of these models were functionally comparable across a wide range of values for $\epsilon$, to the best of our ability. However, previous studies have shown that DP doesn’t make the model learn all classes in a comparable fashion Suriyakumar et al. (2021); Fioretto et al. (2022) as a result, we report the accuracy (acc) as a measure of Perf Comp. and agreement on “hard prediction” to measure Alignment on a class-wise basis, allowing for a more nuanced comparison.
We train the non-private and private models fixing all the hyperparameters except for the number of epochs, as private models need more computation to learn due to the heavy regularization DP introduces in the training Ponomareva et al. (2023). However, to make a fair comparison we have fixed the number of all hyperparameters in all the private models with different $\epsilon$ [^11] We set the no. of epochs as 50 for all private models; it yielded competitive accuracy across model types. However, within $13-16$ epochs, all non-private models achieved accuracy $>95\%$..
However, while training we discovered an issue in the vanilla architecture of the aforementioned networks: they utilize batch normalization (BatchNorm). Nevertheless, BatchNorm normalizes a sample based on the statistics of the batch it is in. This means the same sample can get different normalized values depending on the other samples in the batch. For differential privacy, each sample’s privacy needs to be independently preserved. Since BatchNorm depends on other samples, it violates this principle and leaks information about other samples. Yousefpour et al.  Yousefpour et al. (2022) advises replacing BatchNorm layers with privacy-friendly options like Group Normalization, Layer Normalization, Instance Normalization etc. From the engineering perspective, we have to select one such replacement that scales with sufficiently large datasets without hampering the privacy bounds. Based on previous empirical evidence Subramani et al. (2021), we replace the BatchNorm layers with GroupNorm layers in all non-private models along with their private counterparts, as GroupNorm does not alter the base architecture drastically, scales well and adheres to the privacy principle strictly. We employ the DP-SGD algorithm for the private training Abadi et al. (2016). The non-private models’ accuracies are reported in Table 1 and in Figure 1, 2, 3 for private counterparts.
@@
Figure 1: Performance of the post-hoc explainers (DenseNet-121)
Figure 2: Performance of the post-hoc explainers (ResNet-34)
Figure 3: Performance of the post-hoc explainers (EfficientNet-v2)

### v4

@@ 4 Localization Assumption & Privacy Invariance Score
Right-to-Privacy (RTP) (Thomson, 1975) and Right-to-Explanation (RTE) (Vredenburgh, 2022) are two inalienable aspects of modern-day ML software. We have already mentioned that previous works have shown potential security breaches leveraging the explanation in section 3.1. So, RTE can hamper RTP but is the converse true? If yes, how can we quantify the same? We first propose the desiderata for explanations in this setting and then run extensive experiments to quantitatively judge the explainers.
@@ 4.1 Description of the setting
It is shown that adversaries can leverage sensitive information from explanations but due to post-processing of DP, any DP model will always output a private explanation (Dwork et al., 2014). We want to check the extent to which the explanations from a DP model can be used as a proxy for that of the non-private model here. Formally, for a non-private model $\mathcal{M}$, its private counterpart $\mathcal{M^{\prime}}$; consider a set of post-hoc explainers $\mathcal{E}$ used for auditing $\mathcal{M}$ and $\mathcal{M^{\prime}}$. For an input $x\in X$ (where $X$ is the valid input space), each $\mathcal{I}\in\mathcal{E}$ produces explanations $s$ and $s^{\prime}$ of $x$ for $\mathcal{M}(x)$ and $\mathcal{M^{\prime}}(x)$ respectively, Privacy Invariance Score (PIS) over a given tuple $(\mathcal{M},\mathcal{M^{\prime}},x,\mathcal{I})$ is defined as follows.
Definition 1. Privacy Invariance Score (PIS): Given a tuple $(\mathcal{M},\mathcal{M^{\prime}},x,\mathcal{I})$, given $\mathcal{M}(x)$ = $\mathcal{M^{\prime}}(x)$, PIS is defined as $sim(s,s^{\prime})$, where $sim(\cdot,\cdot)$ is a similarity measure, and $s=\mathcal{I}(\mathcal{M},x)$ and $s^{\prime}=\mathcal{I}(\mathcal{M^{\prime}},x)$.
@@ 4.2 Localization Assumption
As discussed earlier, different works evaluated explainers in various ways. These ad-hoc norms are often unique to each paper and inconsistent (Jacovi and Goldberg, 2020). However, practitioners have proposed some necessary axiomatic desiderata (Lyu et al., 2024) such as the well-established Implementation Invariance (II) criterion proposed by Sundararajan et al. (Sundararajan et al., 2017).
According to II if $\mathcal{I}$ is faithful and $\mathcal{M}(x)=\mathcal{M^{\prime}}(x)$ $\forall x\in X$, i.e. $\mathcal{M,M^{\prime}}$ are functionally equivalent (FE) then $s=s^{\prime}$ $\forall x$. Commonly used explainers like Integrated Gradient (Sundararajan et al., 2017), SmoothGrad (Smilkov et al., 2017), DeepLift (Li et al., 2021), layerwise relevance propagation (LRP) (Montavon et al., 2019) etc had been extensively accessed based on II (Sundararajan et al., 2017).
Following these lines of prior explainability research, Jacovi et al. later formally proposed The Model Assumption (MA) (Jacovi and Goldberg, 2020):
@@
Nevertheless, in our setting, as the accuracy of $\mathcal{M^{\prime}}$ is generally less (Abadi et al., 2016), and we cannot say $\mathcal{M}$ and $\mathcal{M^{\prime}}$ are FE (Sundararajan et al., 2017). Furthermore, as $X$ in practice can be arbitrarily large (theoretically could be countably infinite), quantifying whether an explainer ($\mathcal{I}$) is sufficiently trustworthy or not is often impractical. Also, finding a counterexample that violates the condition implying $\mathcal{I}$ is not faithful is computationally expensive. Hence, we modify MA and propose the localization assumption for evaluating the explainers’ quality in our setting.
@@
    Alignment with the original model (Alignment): The proxy model must closely align with the original model in its predictions. In our study, we measure the same with the agreement on the “hard predictions" obtained from $\mathcal{M}$ and $\mathcal{M^{\prime}}$ over the test set. (Dis)agreement has been extensively used in previous studies due to its its simplicity and interpretability (Klabunde et al., 2024).
@@ 4.3 The Notion of Similarity
As previously discussed in PIS, we aim to measure the similarity between pairs of explanations; it is essential to account for two key factors here: the context of comparison and human understanding of that comparison. In this framework, given $(\mathcal{M},\mathcal{M^{\prime}},x,\mathcal{I})$ and $\mathcal{M}(x)=\mathcal{M^{\prime}}(x)$, we seek to investigate two primary aspects:
@@
Firstly, explanations typically include both positive and negative attributions [^4] Saliency, for example, outputs only positive attributions as absolute values., where positive attributions favor the classification, and negative ones do not. To evaluate explanations’ similarity, we first compute a disagreement score (DS, measured in %), by measuring the mismatch of corresponding attribution type (negative or positive) in ($s,s^{\prime}$). DS is a (crude) measure for assessing how much the reasoning of the models, as indicated by $\mathcal{I}$, diverges at the pixel or element level. In several highstakes (e.g. biomedical), as both position and attribution are important, we employ DS primarily as a sanity check in our case study with a threshold (hyperparameter) of 15%.
Secondly, after measuring DS and eliminating the explainers not adhering the threshold, we assess the attribution correlation for the rest. We start by segregating the positive attributions from the negative ones, as the former contribute positively to the classification, while the later may (partially or completely) contribute to other classes or contribute negatively to the output class or a combination of these two (Ancona et al., 2017; Samek et al., 2017). Consequently, we compute the Kendall’s tau between pairwise positive pixels of s and $s^{\prime}$ to measure PIS. While alternative notions for similarity like cosine similarity or $L_{P}$ norms exist, cosine is unsuitable in high dimensions due to near-orthogonality of random vectors (Mohammadi and Petridis, 2022; Wyner, 1967), and $L_{P}$ norms require normalization, losing one degree of freedom unlike correlation.
@@ 5.1 Description of the Dataset
Our dataset comprises 2,000 Pneumonia cases sourced from the Chest X-ray dataset by (Patel, 2020), and 2,000 TB cases randomly sampled from the NIAID TB Portal Program dataset (National Institute of Allergy and Infectious Diseases, ). To create the ‘Normal’ subset, we include an equal split of 1,000 unaffected Pneumonia samples from the unaffected class in the aforementioned Chest X-ray dataset (Patel, 2020) and 1,000 unaffected TB samples from Rahman et al. (Rahman et al., 2020), totaling 2,000 normal cases. For evaluation, the test set contains 200 images for each class.
@@
This paper, unlike prior studies, is not focused on the empirical studies of private models in healthcare (Naresh et al., 2023; Khalid et al., 2023). Instead, we concentrate on the applicability of commonly used explainers in a privacy-preserving environment. Our aim is to investigate whether RTE and RTP can be achieved simultaneously. As a result, rather than a large dataset where extensive experimentation on several types of CNNs with DP under a closely controlled environment is severely challenging and computationally demanding (Ponomareva et al., 2023), we focus on first making a dataset of an appropriate size where we have considered the commonly available and utilized disease class from the previous studies. However, after our primary experiment, we, within the limits of feasibility, experimented with another benchmark dataset: CIFAR-10, and we arrived at a similar conclusion. Details can be found in Appendix A.
5.2 Choosing the Privacy Budget ($\epsilon$)
Selecting $\epsilon$ is critical to balancing privacy requirements and utility. The theoretical school advocates for small $\epsilon$s (e.g. $\epsilon\leq 1$), offering strong worst-case guarantees (for e.g. $\epsilon=0.1$ theoretically limits MIA success rates to 52.5%). In contrast, industrial settings often adopt large $\epsilon$ (e.g. $\epsilon>7$), which, while practical, theoretically yield MIA success rate exceeding 99%.
This dichotomy arises from differing assumptions (Lowy et al., 2024): the theoretical school presumes adversaries with near-complete knowledge of training data and uniform privacy guarantees for all datapoints. However, recently Lowy et al. (Lowy et al., 2024) highlights the impracticality of these assumptions and showed that even $\epsilon\geq 7$ can provide adequate defense against existing MIAs.
While this study does not propose guidelines for choosing $\epsilon$ in biomedical applications, we adopt Google’s three-tier framework (Ponomareva et al., 2023):
Tier 1: $\epsilon\leq 1$: strong theoretical guarantees, low utility;
Tier 2: $\epsilon\leq 10$: practical trade-offs, google advocates this tier primarily;
Tier 3: $\epsilon>10$: high vulnerability, thus excluded from this study.
@@ 5.2 Choosing the Privacy Budget ($\epsilon$)
Our analysis examines $\epsilon\in\{0.4,0.7,1,4,7,10\}$, spanning tier 1 and tier 2, to comprehensively evaluate privacy-utility trade-offs.
@@ 5.3 Description of the Explainers
We primarily selected gradient-based explainers, which are predominantly used in computer vision tasks (Buhrmester et al., 2021) — specifically, Saliency, SmoothGrad, Integrated Gradients, Grad-Shap and Grad-CAM (details are in Appendix B). These gradient-based explainers rely on the sensitivity the model shows to the features subject to the output it obtains. In other words, these explainers leverage the gradient of the output (or any selected class of interest) w.r.t. the input features555it can either directly be the element(s) of the given input or the intermediate ones coming from a layer inside the model.. For example, Saliency computes the partial derivatives of the output w.r.t. the input given, Integrated Gradients computes the average gradient while the input varies along a linear path from a baseline. SmoothGrad involves adding noise to the input image and generating multiple saliency maps, each corresponding to a slightly different noisy version of the input image; the saliency maps are then averaged to produce a ‘smoothed’ saliency map. Grad-CAM also involves taking gradients of the target output with respect to the given layer. From perturbation-based explainers, we included only SHAP (Lundberg and Lee, 2017), specifically Grad-Shap. Grad-Shap, in a sense, can be viewed as an approximation of Integrated Gradients by computing the expectations of gradients for different baselines. Core perturbation methods are less widely adopted for computer vision, and they tend to disobey the “data manifold hypothesis” (Fefferman et al., 2016), as described in the next paragraph. Furthermore, perturbation-based methods like LIME (Ribeiro et al., 2016) assign weights to super-pixels rather than individual pixels. As a result, we cannot directly compare the same with other methods we chose, which output per-pixel or per-element attribution values, Occlusion can lead to OOD artifacts and is not widely used (Chang et al., 2018). Additionally, some other explainers, such as DeepLIFT, LRP disregard even basic faithfulness tests (Sundararajan et al., 2017).
We did not consider model-agnostic perturbation-based explainers (e.g., KernelSHAP, LIME, Occlusion) due to their relatively limited adaptation in computer vision tasks and several criticisms. Firstly, these are known to disobey the “data manifold hypothesis” (Fefferman et al., 2016): Kumar et al. (Kumar et al., 2020) and Sundararajan et al. (Sundararajan and Najmi, 2020) showed these methods often operate outside the data’s true distribution. Furthermore, Shokri et al. (Shokri et al., 2021) acknowledged the fact that “little is known on how models generally perform outside of the data manifold. In fact, it is not even clear how one would measure performance of a model on points outside of the training data distribution: they do not have any natural labels… ”. To the best of our knowledge, not only vanilla models but also how their DP counterparts behave outside the data manifold is unknown. In fact, how different the manifold of DP models is, compared to its non-DP counterpart, and how these perturbation-based explainers may behave subject to DP models’ decision boundaries and manifolds, has not been extensively studied prior. Therefore, whether, in an ad-hoc manner, we can even employ these explainers to DP models or the suitability of these explainers in the context of DP models is not clear to us, and we’d like to leave this as an open research question to the community. However, as mentioned earlier, we considered Grad-Shap as it closely resembles to Integrated Gradients.
@@ 5.4 Description of the Networks
Disease detection using chest X-ray is a well-regarded case study. In previous works (Al-qaness et al., 2024), the choice of CNNs is significantly prominent, and in this work, we have also chosen the famous CNNs that have been used in previous research. Particularly, We selected ResNet-34 (He et al., 2016), EfficientNet-v2 (‘small’ version) (Tan and Le, 2021), and DenseNet-121 (Huang et al., 2017) for our analysis due to their competitive performance. Importantly, we ensured that the DP counterparts of these models were functionally comparable across a wide range of values for $\epsilon$, to the best of our ability. However, previous studies have shown that DP doesn’t make the model learn all classes in a comparable fashion (Suriyakumar et al., 2021; Fioretto et al., 2022) as a result, we report $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$ as a measure of Perf Comp. and agreement on “hard prediction” to measure Alignment on a class-wise basis, allowing for a more nuanced comparison. We detailed our experimental setup in Appendix C.
@@
Figure 1: Performance of explainers (ResNet-34)
Figure 2: Performance of explainers (DenseNet-121)
Figure 3: Performance of explainers (EfficientNet-V2)